# Analyzing Character Representation in Media Content using Multimodal Foundation Model: Effectiveness and Trust

### v1

@@ 4.4. Discussion
Overall, our character representation and visualization tool for media content presented the information at a satisfactory level. Most participants were able to understand the presented information for gender, age and their intersection (Q2.2-3,5-6) and the model-related information (Q2.7-8) quite accurately. There is quite some uncertainty around people’s understanding of what the percentages represented (Q2.1), which was also raised by 3 participants in Q3.10. The accuracy of responses in Q2.4 is also uncertain but this might be due to the broad nature of the question.
@@
The participants’ liking of using our tool was moderate (Q3.9), while the overall rating of our tool (Q3.8) was above average. Based on all this data we believe that our tool is still relevant, interesting, and useful. 40% of the participants think that the existing streaming platforms/movie database sites are slightly or not at all informative about the diversity/inclusivity of characters in films and tv series (Q1.6), while some identified a broader applicability of such tools not only in film industry, but in media more broadly, and beyond. Nonetheless, based on the data we collected, there is scope for
@@
  • including more demographics, like ethnicity/race, and more contextual analyses of the presented characters,
@@
We believe that AI and visualization are valuable tools for analyzing the character representation in media and communicating these analyses to the broader audience. In the era of Over-the-top production of media content, tools that could automate such analyses would help people better understand the offered content and identify what is relevant and interesting to them through a perspective not currently available to them. But there is need for more work to better map the audience’s expectations from such tools, and strike a balance in the complex interplay of explainability, accuracy, and trust (Papenmeier et al., 2022).
@@ 5. Conclusion
Recent advances in AI has made it possible to analyze media content with reasonable accuracy at scale and generate character representation analytics. Getting the audience in the loop within this process is important. We need to understand what matters to them and whether they would trust such an analysis system. We built an open-source tool based on the CLIP model to analyze the character representation across dimensions of age and gender and visualize the extracted analytics. We conducted a user study to provide empirical evidence on the usefulness and trustworthiness of the AI-generated results for three full length movies presented in form of our visualizations. Participants understood well the presented information and deemed the tool overall useful but there is a request for more tailored visualizations that include more demographic categories and contextual analyses of the character portrayals. Participants did not reject the use of AI in this context, they found the presented information about the AI confidence and bias very useful, but their trust in our gender and age models was mediocre and low, respectively. These are just some initial findings but more work is required towards this direction. In the era of over-the-top production of media content, tools like this one will become more and more important in helping people identify content that is relevant and interesting to them.
@@
The benchmarking of the models was based on a free and publicly available dataset. We legally acquired the video content we analyzed. The user study was approved by relevant authority at the authors’ institution. All participants data collected is anonymized and therefore individuals cannot be personally identified. We carefully considered the risks for participants in taking part in the user study. We minimized the risk of participant fatigue by setting a maximum time limit of 30 minutes. Our tool should be only used with legally acquired media content.
@@ Acknowledgment
Hidden for anonymity.

### v2

@@ 4.4. Discussion
Summary of findings. Our character representation and visualization tool for media content presented the information at a satisfactory level. Most participants were able to understand the presented information for gender, age and their intersection (Q2.2-3,5-6) and the model-related information (Q2.7-8) quite accurately. There is quite some uncertainty around people’s understanding of what the percentages represented (Q2.1), which was also raised by 3 participants in Q3.10. The accuracy of responses in Q2.4 is also uncertain but this might be due to the broad nature of the question.
@@
The participants’ liking of using our tool was moderate (Q3.9), while the overall rating of our tool (Q3.8) was above average. These findings seem to suggest that our tool is still relevant, interesting, and useful. 40% of the participants think that the existing streaming platforms/movie database sites are slightly or not at all informative about the diversity/inclusivity of characters in films and tv series (Q1.6), while some identified a broader applicability of such tools in media more broadly, and beyond (Q3.11). Nonetheless, based on the data we collected, there is scope for
@@
  • including more demographics, like ethnicity/race, and more contextual analyzes of the presented characters,
@@
Future directions. Our tool’s development is characterized by modularity and input standardisation, which make it easily extendable to include e.g., other demographic dimensions, more context or interactivity. Due to the lack of previous studies in this topic, we kept the design of the study as simple as possible to avoid overwhelming people and allowing space to ask about various other aspects of the tool than its design, such as trust. Comparative studies are required to test the effects of alternative (refined) designs of the visualization or the effects of AI- vs. human-generated analytics. While we started with a simple design and study, those could be future steps in this work.
We believe that AI and visualization are valuable tools for analyzing the character representation in media and communicating these analyzes to the broader audience. In the era of Over-the-top production of media content, tools that could automate such analyzes would help people better understand the offered content and identify what is relevant and interesting to them through a perspective not currently available to them. But there is need for more work to better map the audience’s expectations from such tools, and strike a balance in the complex interplay of explainability, accuracy, and trust (Papenmeier et al., 2022).
@@ 5. Conclusion
Recent advances in AI has made it possible to analyze media content with reasonable accuracy at scale and generate character representation analytics. Getting the audience in the loop within this process is important. We need to understand what matters to them and whether they would trust such an analysis system. We built an open-source tool, which is not an off-the-shelf AI facial recognition system but is a frontier multimodal model employed in a novel way: based on the CLIP model we analyze the character representation across dimensions of age and gender and visualize the extracted analytics. We conducted a user study to provide empirical evidence on the usefulness and trustworthiness of the AI-generated results for three full length movies presented in form of our visualizations. Participants understood well the presented information and deemed the tool overall useful but there is a request for more tailored visualizations that include more demographic categories and contextual analyzes of the character portrayals. Participants did not reject the use of AI in this context, they found the presented information about the AI confidence and bias very useful, but their trust in our gender and age models was mediocre and low, respectively. These are just some initial findings but more work is required towards this direction. In the era of over-the-top production of media content, tools like this one will become more and more important in helping people understand the film through the lens of film consumption and decision-making about cultural or social content.
@@
The benchmarking of the models was performed on a free and publicly available dataset. We legally acquired the video content we analyzed. The user study was approved by relevant authority at the authors’ institution. All participants data collected is anonymized and therefore individuals cannot be personally identified. We carefully considered the risks for participants in taking part in the user study. We minimized the risk of participant fatigue by setting a maximum time limit of 30 minutes. Our tool should be only used with legally acquired media content. The intellectual property of creatives, fairness and consent are key issues our research is committed to.
@@ Acknowledgment
This work was funded by the Leverhulme Trust (RPG-2023-091).